# Synthetic Prior for Few-Shot Drivable Head Avatar Inversion

### v1

@@ Abstract
We present SynShot, a novel method for the few-shot inversion of a drivable head avatar based on a synthetic prior. We tackle two major challenges. First, training a controllable 3D generative network requires a large number of diverse sequences, for which pairs of images and high-quality tracked meshes are not always available. Second, state-of-the-art monocular avatar models struggle to generalize to new views and expressions, lacking a strong prior and often overfitting to a specific viewpoint distribution. Inspired by machine learning models trained solely on synthetic data, we propose a method that learns a prior model from a large dataset of synthetic heads with diverse identities, expressions, and viewpoints. With few input images, SynShot fine-tunes the pretrained synthetic prior to bridge the domain gap, modeling a photorealistic head avatar that generalizes to novel expressions and viewpoints. We model the head avatar using 3D Gaussian splatting and a convolutional encoder-decoder that outputs Gaussian parameters in UV texture space. To account for the different modeling complexities over parts of the head (e.g., skin vs hair), we embed the prior with explicit control for upsampling the number of per-part primitives. Compared to SOTA monocular methods that require thousands of real training images, SynShot significantly improves novel view and expression synthesis.
Figure 1: Given a few input images (left), SynShot generates a personalized 3D Gaussian avatar that renders from new viewpoints and unseen expressions (right). To compensate for the missing information in the input images, we leverage a generative Gaussian head avatar trained on a diverse synthetic head dataset as a 3D prior. To generate a personalized avatar, SynShot uses pivotal tuning to project the input onto the prior’s manifold, then fine-tunes the appearance networks to fit the input images. While being applicable to an arbitrary number of input images, the figure shows novel views and expressions (right) for an avatar generated from only three input images (left).
footnotetext: Work done while Wojciech Zielonka was interning at Google in Zurich, Switzerland
@@ 1 Introduction
Another recent body of work addresses the problem of building 3D head avatars from one or few input images, [73, 8, 11]. However, their rendering quality and fidelity are typically lower than those of methods trained on large datasets (e.g., [47, 82]). To improve quality, some methods [76, 71, 72] first learn a multi-identity head model that is used as prior when optimizing for the personalized avatar. Training these head priors requires a large-scale multi-view image dataset that is expensive and time-consuming to capture. The expressive power of the prior is strongly influenced by several factors: the diversity of the training data (e.g., ethnicity, age, facial features, expressions), the multi-view capture hardware setup (i.e., lighting, view-density, calibration quality, frame-rate), and the quality of the data pre-processing (e.g., mesh tracking, background masking).
In contrast to the previous work that focuses on expensive real data, we overcome these limitations and propose SynShot, a new method that builds a prior on synthetic data, and adapts to a real test subject requiring only a few input images. Building on the success of ML models trained on synthetic data for tasks like 3D face regression [54], 2D landmark prediction [66], rigid face alignment [3], and few-shot head reconstruction [64, 74, 6], SynShot is trained solely on a large synthetic dataset generated from 3DMM samples and diverse assets. Synthetic data offers complete control over dataset creation to meet size and diversity needs for training an expressive head prior, eliminating the need for costly capture hardware and addressing privacy concerns with real subjects. The benefits brought by synthetic data come at the cost of having to handle the domain gap between the trained head prior and real images captured “in the wild”. To effectively bridge this gap, we first fit the synthetic prior to real images and then fine-tune the prior weights to the real data using the pivotal tuning strategy proposed in [49]. With as few as three input images, SynShot reconstructs a photorealistic head avatar that generalizes to novel expressions and camera viewpoints (Fig. 1). Our results show that our method outperforms state-of-the-art personalized monocular methods [80, 67, 56] trained on thousands of images each. Our method represents head avatars using 3D Gaussian primitives [28], where Gaussian parameters are generated by a convolutional architecture in the UV (texture) plane, similar to related work [39, 82, 52].
@@
    A novel generative method based on a convolutional encoder-decoder architecture that is trained on extensive synthetic data to produce controllable 3D head avatars.
@@ 2 Related Work
Figure 2: Pipeline overview. Given an extracted texture in a neutral position, rasterized position map, and PCA expression deltas our network regresses Gaussian maps which later undergo 3DGS splatting.
@@
3D Morphable Models (3DMM) [4, 45, 38, 14] have long been used for creating facial avatars. When paired with generative models for textures [21, 35, 36, 41], 3DMMs can be optimized from in-the-wild images. Techniques such as inverse rendering [13], diffusion-based inpainting [44], and pivotal-tuning [49, 37] are used to disentangle appearance from identity. Neural radiance fields (NeRF) [43] and 3D Gaussian representations (3DGS) [28] have also been widely used for avatar reconstruction. EG3D [7] employs features on tri-planes, enabling consistent 3D face generation and inversion from in-the-wild images. PanoHead [2] extends EG3D through tri-grids to achieve full 360-degree generation of static human heads. Gaussian3Diff [34] further improves quality by replacing neural features with 3D Gaussians. Rodin [64] and RodinHD [74] leverage an extensive dataset of synthetic humans to train a triplane-based avatar generator used to invert in-the-wild images; however, the results remain confined to the synthetic domain and avatars are not drivable.
Diner [46] incorporates depth information, while Preface [5] trains a volumetric prior on synthetic human data and fits it to a few input images to match a subject’s likeness. Cafca [6] extends Preface to better generalize to static but arbitrary facial expressions. In contrast, our method not only bridges the domain gap from synthetic to real but also produces animatable avatars. MofaNeRF [79] learns a NeRF embedded on a parametric model and, as NeRFace [18], conditions the NeRF on expression codes, with additional shape information for generalization. Another NeRF-based parametric model is HeadNeRF [25]. Ultimately, these methods replace traditional textures models with neural volumetric representations, to better capture features such as hair. As NeRFs, they suffer from slow rendering times and low-resolution outputs.
Recent methods by Xu et al. [72] are conceptually similar to HeadNeRF and MofaNeRF; however, instead of embedding NeRF [43] on a mesh, they employ 3DGS [28]. GPHM [72] uses a series of MLPs to generate Gaussian primitives attached to a parametric model, enabling expression control and inversion, though it conditions only the avatar’s shape. In contrast, our approach incorporates both shape and color information. The follow-up work GPHMv2 [71] extends GPHM with a dynamic module for improved reenactment control and a larger dataset, further enhancing quality. HeadGAP [76] also models avatars using MLPs, utilizing part-based features and additional color conditioning to improve quality. While these methods embed primitives directly on the mesh surface, our approach explicitly learns the primitive parameters by modeling their distribution via a VQ-VAE, eliminating the need for a guiding mesh during the test time as the shape is captured within our latent space.
@@ 3.1 Preliminaries
Using a 3DMM, we represent a base 3D mesh as $\mathbf{S}=\bar{\mathbf{S}}+\boldsymbol{\delta}\mathbf{B}_{id}+\boldsymbol{\gamma}\mathbf{B}_{expr}$, where $\bar{\mathbf{S}}$ is the average shape, $\mathbf{B}_{(expr,id)}$ are the bases for expression and identity, and $\boldsymbol{\delta}$, $\boldsymbol{\gamma}$ denote the corresponding coefficients. Additionally, we use linear blend skinning (LBS) for head rotation around the neck with pose corrective offsets, and to rotate the eyeballs.
@@ 3.2 Gaussian Prior Model
We define the primitive positions in the 3DMM space using only shape and expression. Global rotation, translation, and linear blend skinning (LBS) are factored out and applied to the primitives just before splatting to place them in the correct world space. We compute initial per-part Gaussian parameters for our primitives. Note that we do not use a fixed canonical space [31, 82, 22], as our initialization is derived from predictions. We first obtain positions by sampling $\boldsymbol{\phi}_{r}=\mathcal{B}(\boldsymbol{\phi},u_{r},v_{r})$, for $r\in\{face,hair\}$. Next, for each $\boldsymbol{\phi}_{r}$, we compute nearest neighbor distance and initialize scale as $\boldsymbol{\sigma}_{r}=\min_{j\neq i}\|\boldsymbol{\phi}_{r_{i}}-\boldsymbol{\phi}_{r_{j}}\|_{2}$. Initial opacity is set to $\boldsymbol{\alpha}=0.7$. Finally, the per-part rotations are computed as $\boldsymbol{\theta}_{r}=\begin{bmatrix}\frac{\mathbf{T}}{\|\mathbf{T}\|}&\frac{\mathbf{B}}{\|\mathbf{B}\|}&\frac{\mathbf{N}}{\|\mathbf{N}\|}\end{bmatrix}\in\mathbb{R}^{h\times w\times 3\times 3}$, based on the image space gradient:
@@
Additionally, the position maps $\hat{\boldsymbol{x}}_{\mathrm{verts}}$ and expression maps $\hat{\boldsymbol{x}}_{\mathrm{expr}}$ are supervised by $\mathcal{L}_{\text{geom}}=\delta\mathcal{L}_{L1}$. The final loss is defined as $\mathcal{L}=\mathcal{L}_{\text{color}}+\mathcal{L}_{\text{geom}}$. Moreover, we apply $L_{2}$ regularization on position, scale, opacity, and the FC part of the spherical harmonics coefficients: The final loss is defined as $\mathcal{L}=\mathcal{L}_{photo}+\mathcal{L}_{reg}$. We train our network end-to-end using 8 GPUs Nvidia H100 with batch size 16 (2 per GPU). We optimize the network for 500K iterations with the Adam optimizer [29] with lr=$1.3e^{-5}$ and multi-step scheduler which decays the learning rate every milestone by gamma=0.66.
@@ 3.3 Few-shot Avatar Reconstruction
To bridge the gap between real (ITW) and synthetic avatars, we carefully designed a two-stage inversion process based on pivotal fine-tuning [49].
@@
First, we optimize the encoder $\boldsymbol{E}_{\mathrm{id}}$ while keeping the rest of the network fixed such that we recover $\boldsymbol{z}_{\mathrm{id}}$. Note that $\boldsymbol{E}_{expr}$ remains unchanged as it should model independent expressions. Once $\boldsymbol{E}_{id}$ is fine-tuned, we fix its predicted identity latent code $\boldsymbol{z}_{\mathrm{id}}$, we fine-tune the decoders $\{\boldsymbol{D}_{\mathrm{feat}},\boldsymbol{D}_{\mathrm{id}},\boldsymbol{D}_{\mathrm{expr}}\}$ and the regressors $\{\boldsymbol{R}_{color},\boldsymbol{R}_{gauss}\}$ for the hair and face regions (Figure 4).
@@ 3.4 Synthetic Dataset
Figure 6: Cross-reenactment comparison of SynShot inversion using only 3 views to state-of-the-art (SOTA) methods: INSTA [80], Flash Avatar (FA) [67], and Splatting Avatar (SA) [56], each of which was trained on an average of 3000 frames. It is evident that, without a strong prior, these methods lack generalization to novel expressions and views.
@@ 4 Results
Refer to captionOurs (3)FA (3)SA (3)INSTA (3)Ours (987)FA (987)SA (987)INSTA (987)
Figure 7: Novel view evaluation reveals severe overfitting in personalized methods such as INSTA [80], Flash Avatar (FA) [67], and Splatting Avatar (SA) [56]. Each trained on all available (987) or three frames (3). In contrast, our method, utilizing only three (3), demonstrates significantly better generalization, benefiting from our synthetic prior.
Refer to captionSourceOursHN [25]MN [79]PH [2]
Figure 8: Single-shot inversion methods compared to SynShot. HeadNeRF [25] (HD), MofaNeRF [79] (MF) and PanoHead [2] (PH). Using only one image, our method maintains geometric consistency, matching or exceeding other state-of-the-art approaches.
@@
For inversion-based methods, we compare PanoHead [2], HeadNeRF [25], and MofaNeRF [79], which project the input image onto their latent space and refine the network. Figure 8 demonstrates that the results of HeadNeRF and MofaNeRF are blurry and lack person-specific details. While PanoHead provides faithful reconstructions, their result is not animatable. Moreover, Figure 9 presents a novel view evaluation of challenging long hair and beard inversion, demonstrating strong generalization capabilities of SynShot. Additional results can be found in the supplementary material. To evaluate the performance of SynShot without introducing bias, we selected training frames from $\{F_{n}\}_{n=1}^{16}=\{1,\dots,987\}$, where $F_{n}$ denotes the Fibonacci sequence. We then use progressive farthest point sampling in the 3DMM expression space to sample a given number of frames from the training set.
@@
Our combined dataset consists of eleven monocular sequences ($512\times 512$ resolution), many of which are in-the-wild videos with very limited head motion, resulting in a low error as the test sequences closely resemble the training data, leaving limited room to assess diversity. To address this and accurately measure the effective error, we trained each method on a varying number of frames (Figure 11), corresponding to frames used in our inversion pipeline. The reconstruction error is evaluated on 600 test frames. Figure 10 demonstrates the effectiveness of our inversion, particularly with up to 233 training frames. Due to the lack of a strong prior, monocular methods fail in low training frame regimes, and, even with larger training datasets, they do not perform well and produce artifacts. Please refer to the supp. mat. for qualitative evaluation.
Ground Truth
Figure 11: The quality of our avatars’ inversion improves as we use more training frames and learn out-of-distribution illumination and appearance. Note that the single-frame reconstruction is based on a neutral expression with a closed mouth. See supp. mat. for more results.
@@
Refer to captionSourceOursINSTA [80]SA [56]FA [67]
Figure 12: Cross-reenactment with Limited Frames. Despite achieving good results on test frames using only 13 frames, all monocular baseline methods suffer from heavy artifacts in the expression transfer task. Full evaluation in the supp. mat.
We would like to emphasize the importance of evaluating cross-reenactment, which often reveals issues with generalization and overfitting; however, these aspects are frequently underemphasized, as evaluation sequences are commonly not sufficiently challenging. For instance, Figure 10 indicates that 1̃3 frames may be sufficient for monocular methods to perform well on the test set. However, Figure 12 shows that models trained on this number of frames fail severely on cross-reenactment tasks. In the supp. mat. we present a full evaluation. Therefore, despite achieving high-quality results, most monocular methods [24, 80, 18, 67, 56] struggle with cross-reenactment involving novel expressions and views. Without a strong prior, these methods frequently exhibit artifacts when driven by out-of-distribution sequences. In contrast, our method, leveraging only three images and a synthetic prior with effective shape-expression disentanglement, is able to invert an avatar that significantly outperforms state-of-the-art models trained on thousands of frames. Figure 18 demonstrates cross-reenactment, with the leftmost column serving as the source for expression and view. This shows that incorporating a strong prior enhances the visual quality.
Loss L1 $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
$\mathcal{L}_{photo}+\mathcal{L}_{VGG}$ 0.0244 0.0839 0.9058 23.1191
Table 1: Ablation for our inversion losses.
Single Image Inversion.
To project a single image into a 3D avatar, a strong prior is essential, as this is a highly unconstrained optimization problem. Similar to our approach, HeadNeRF [25] and MofaNeRF [79] produce controllable avatars. However, they lack fine-grained control, and PanoHead [2] reconstructs only static avatars. Moreover, all of these methods are trained on real or combinations of real and synthetic data, which does not introduce a sim-to-real domain gap. We tackle a more challenging problem, aiming to create a photorealistic avatar with fine-grained control based solely on a synthetic prior. Despite this, our single-image reconstructions surpass other methods. Figure 8 shows two subjects from our dataset where the single in-the-wild source frame was inverted. Our results present photorealistic reconstructions with plausible side views, highlighting the importance of the synthetic prior.
VQ-VAE Architecture Ablation.

### v2

@@ Abstract
We present SynShot, a novel method for the few-shot inversion of a drivable head avatar based on a synthetic prior. We tackle three major challenges. First, training a controllable 3D generative network requires a large number of diverse sequences, for which pairs of images and high-quality tracked meshes are not always available. Second, the use of real data is strictly regulated (e.g., under the General Data Protection Regulation, which mandates frequent deletion of models and data to accommodate a situation when participant’s consent is withdrawn). Synthetic data, free from these constraints, is an appealing alternative. Third, state-of-the-art monocular avatar models struggle to generalize to new views and expressions, lacking a strong prior and often overfitting to a specific viewpoint distribution. Inspired by machine learning models trained solely on synthetic data, we propose a method that learns a prior model from a large dataset of synthetic heads with diverse identities, expressions, and viewpoints. With few input images, SynShot fine-tunes the pretrained synthetic prior to bridge the domain gap, modeling a photorealistic head avatar that generalizes to novel expressions and viewpoints. We model the head avatar using 3D Gaussian splatting and a convolutional encoder-decoder that outputs Gaussian parameters in UV texture space. To account for the different modeling complexities over parts of the head (e.g., skin vs hair), we embed the prior with explicit control for upsampling the number of per-part primitives. Compared to SOTA monocular and GAN-based methods, SynShot significantly improves novel view and expression synthesis.
Figure 1: Given a few input images (left), SynShot generates a personalized 3D Gaussian avatar that renders from new viewpoints and unseen expressions (right). To compensate for the missing information in the input images, we leverage a generative Gaussian head avatar trained on a diverse synthetic head dataset as a 3D prior.
footnotetext: Work done while WZ was interning at Google in Zurich, Switzerland
@@ 1 Introduction
Another recent body of work addresses the problem of building 3D head avatars from one or few input images, [77, 8, 11]. However, their rendering quality and fidelity are typically lower than those of methods trained on large datasets (e.g., [50, 88]). To improve quality, some methods [81, 75, 76] first learn a multi-identity head model that is used as prior when optimizing for the personalized avatar. Training these head priors requires a large-scale multi-view image dataset that is expensive and time-consuming to capture. Moreover, managing real data under protection laws like GDPR is cumbersome for experimentation and maintenance, as users must periodically (e.g., every 30 days) delete all dataset derivatives and trained models, allowing dataset participants to be removed from both if needed. Alternatively, the FFHQ dataset [29] may be employed, with 4D GAN-based methods [80, 61, 13] constructing an inversion prior from it. However, these approaches tend to exhibit artifacts during novel view synthesis and struggle with preserving identity. In summary, the expressive power of this prior is strongly influenced by several factors: the diversity of the training data (e.g., ethnicity, age, facial features, expressions), the multi-view capture hardware setup (i.e., lighting, view-density, calibration quality, frame-rate), and the quality of the data pre-processing (e.g., mesh tracking, background masking).
In contrast to the previous work that focuses on expensive and cumbersome real data, we overcome these limitations and propose SynShot, a new method that builds a prior solely on synthetic data and adapts to a real test subject requiring only a few input images. Building on the success of ML models trained on synthetic data for tasks like 3D face regression [57], 2D landmark prediction [70], rigid face alignment [3], and few-shot head reconstruction [68, 78, 6], SynShot is trained solely on a large synthetic dataset generated from 3DMM samples and diverse assets. Synthetic data offers complete control over dataset creation to meet size and diversity needs for training an expressive head prior, eliminating the need for costly capture hardware and addressing privacy concerns with real subjects. The benefits brought by synthetic data come at the cost of having to handle the domain gap between the trained head prior and real images captured “in the wild”. To effectively bridge this gap, we first fit the synthetic prior to real images and then fine-tune the prior weights to the real data using the pivotal tuning strategy proposed in [52]. With as few as three input images, SynShot reconstructs a photorealistic head avatar that generalizes to novel expressions and camera viewpoints (Fig. 1). The results show that our method outperforms state-of-the-art personalized monocular methods [85, 71, 59] trained on thousands of images each. Our method represents head avatars using 3D Gaussian primitives [30], where Gaussian parameters are generated by a convolutional architecture in the UV (texture) plane, similar to related work [41, 88, 55].
@@
    A novel generative method based on a convolutional encoder-decoder architecture that is trained on extensive synthetic data only to produce controllable 3D head avatars.
@@ 2 Related Work
Figure 2: Pipeline overview. Given an extracted texture $\boldsymbol{x}_{\mathrm{tex}}$, rasterized position map $\boldsymbol{x}_{\mathrm{verts}}$, and PCA expression deltas $\boldsymbol{x}_{\mathrm{exp}}$ our network utilized VQ-VAE to jointly optimize for two latent space $\boldsymbol{z}_{\mathrm{expr}}$ and $\boldsymbol{z}_{\mathrm{id}}$. The VQ-VAE decoders predict feature map $\boldsymbol{D}_{\mathrm{feat}}(\mathbf{q}(\boldsymbol{z}_{\mathrm{id}}),\mathbf{q}(\boldsymbol{z}_{\mathrm{expr}}))\rightarrow\hat{\boldsymbol{x}}_{\mathrm{feat}}$, identity and color maps $\boldsymbol{D}_{\mathrm{id}}(\mathbf{q}(\boldsymbol{z}_{\mathrm{id}}))\rightarrow\{\hat{\boldsymbol{x}}_{\mathrm{tex}},\hat{\boldsymbol{x}}_{\mathrm{verts}}\}$, and expression deltas $\boldsymbol{D}_{\mathrm{expr}}(\mathbf{q}(\boldsymbol{z}_{\mathrm{expr}}))\rightarrow\hat{\boldsymbol{x}}_{\mathrm{expr}}$. Finally, bilinearly sampled maps are passed to per-part regressors $\boldsymbol{R}_{color}$ and $\boldsymbol{R}_{gauss}$ to obtain primitives to rasterize.
@@
3D Morphable Models (3DMM) [4, 47, 40, 15] have long been used for creating facial avatars. When paired with generative models for textures [22, 37, 38, 43], 3DMMs can be optimized from in-the-wild images. Techniques such as inverse rendering [14], diffusion-based inpainting [46], and pivotal-tuning [52, 39] are used to disentangle appearance from identity. Neural radiance fields (NeRF) [45] and 3D Gaussian representations (3DGS) [30] have also been widely used for avatar reconstruction. EG3D [7] employs features on tri-planes, enabling consistent 3D face generation and inversion from in-the-wild images. PanoHead [2] extends EG3D through tri-grids to achieve a full 360-degree generation of static human heads. Gaussian3Diff [36] further improves quality by replacing neural features with 3D Gaussians. Rodin [68] and RodinHD [78] leverage an extensive dataset of synthetic humans to train a triplane-based avatar generator used to invert in-the-wild images; however, the results remain confined to the synthetic domain and avatars are not drivable.
Diner [48] incorporates depth information, while Preface [5] trains a volumetric prior on synthetic human data and fits it to a few input images to match a subject’s likeness. Cafca [6] extends Preface to better generalize to static but arbitrary facial expressions. In contrast, our method not only bridges the domain gap from synthetic to real but also produces animatable avatars. MofaNeRF [84] and NeRFace [19] condition NeRFs on expression (and shape) codes, while HeadNeRF [26] similarly embeds NeRFs into parametric models. Portrait4D [13] introduces one-shot 4D head synthesis using a transformer-based animatable triplane reconstructor built on the EG3D [7]. Next3D [61] employs GAN-based neural textures embedded on a parametric mesh; however, it suffers from inversion problems. InvertAvatar [80] tackles the shortcomings of Next3D and further refines avatar inversion using few-shot images. Despite relatively good frontal performance, these GAN-based methods often exhibit artifacts, such as identity changes, in novel view synthesis.
Recent methods by Xu et al. [76] are conceptually similar to HeadNeRF and MofaNeRF; however, instead of embedding NeRF [45] on a mesh, they employ 3DGS [30]. GPHM [76] uses a series of MLPs to generate Gaussian primitives attached to a parametric model, enabling expression control and inversion, though it conditions only the avatar’s shape. In contrast, our approach incorporates both shape and color information. The follow-up work GPHMv2 [75] extends GPHM with a dynamic module for improved reenactment control and a larger dataset, further enhancing quality. HeadGAP [81] also models avatars using MLPs, utilizing part-based features and additional color conditioning to improve quality. While these methods embed primitives directly on the mesh surface, our approach explicitly learns the primitive parameters by modeling their distribution via a VQ-VAE [66], eliminating the need for a guiding mesh during the test time as the shape is captured within our latent space.
@@ 3.1 Preliminaries
Using a 3DMM, we represent a base 3D mesh as $\mathbf{S}=\bar{\mathbf{S}}+\boldsymbol{\delta}\mathbf{B}_{id}+\boldsymbol{\gamma}\mathbf{B}_{expr}$, where $\bar{\mathbf{S}}$ is the average shape, $\mathbf{B}_{(id,expr)}$ are the bases for identity and expression, and $\boldsymbol{\delta}$, $\boldsymbol{\gamma}$ denote the corresponding coefficients. Additionally, we use linear blend skinning (LBS) for head rotation around the neck with pose corrective offsets, and to rotate the eyeballs.
@@ 3.2 Gaussian Prior Model
We define the primitive positions in the 3DMM space using only shape and expression. Global rotation, translation, and linear blend skinning (LBS) are factored out and applied to the primitives just before splatting to place them in the correct world space. We compute initial per-part Gaussian parameters for our primitives. Note that we do not use a fixed canonical space [33, 88, 23], as our initialization is derived from predicted position maps. We first obtain positions by sampling $\boldsymbol{\phi}_{r}=\mathcal{B}(\boldsymbol{\phi},u_{r},v_{r})$, for $r\in\{face,hair\}$. Next, for each $\boldsymbol{\phi}_{r}$, we compute nearest neighbor distance and initialize scale as $\boldsymbol{\sigma}_{r}=\min_{j\neq i}\|\boldsymbol{\phi}_{r_{i}}-\boldsymbol{\phi}_{r_{j}}\|_{2}$. Initial opacity is set to $\boldsymbol{\alpha}=0.7$. Finally, the per-part rotations are computed as $\boldsymbol{\theta}_{r}=\begin{bmatrix}\frac{\mathbf{T}}{\|\mathbf{T}\|}&\frac{\mathbf{B}}{\|\mathbf{B}\|}&\frac{\mathbf{N}}{\|\mathbf{N}\|}\end{bmatrix}\in\mathbb{R}^{h\times w\times 3\times 3}$, based on the image space gradient:
@@
Additionally, the position maps $\hat{\boldsymbol{x}}_{\mathrm{verts}}$ and expression maps $\hat{\boldsymbol{x}}_{\mathrm{expr}}$ are supervised by $\mathcal{L}_{\text{geom}}=\delta\mathcal{L}_{L1}$. The final loss is defined as $\mathcal{L}=\mathcal{L}_{\text{color}}+\mathcal{L}_{\text{geom}}$. Moreover, we apply $L_{2}$ regularization on position, scale, opacity, and the FC ($l>=1$) part of the spherical harmonics coefficients: The final loss is defined as $\mathcal{L}=\mathcal{L}_{photo}+\mathcal{L}_{reg}$. We train our network end-to-end using 8 GPUs Nvidia H100 with batch size 16 (2 per GPU). We optimize the network for 500K iterations with the Adam optimizer [31] with lr=$1.3e^{-5}$ and multi-step scheduler which decays the learning rate every milestone by gamma=0.66.
@@ 3.3 Few-shot Avatar Reconstruction
To bridge the gap between in-the-wild and synthetic avatars, we carefully designed a two-stage inversion process based on pivotal fine-tuning [52]. First, we optimize the encoder $\boldsymbol{E}_{\mathrm{id}}$ while keeping the rest of the network fixed such that we recover $\boldsymbol{z}_{\mathrm{id}}$. Note that $\boldsymbol{E}_{expr}$ remains unchanged as it should model independent expressions. Once $\boldsymbol{E}_{id}$ is fine-tuned, we fix its predicted identity latent code $\boldsymbol{z}_{\mathrm{id}}$, we fine-tune the decoders $\{\boldsymbol{D}_{\mathrm{feat}},\boldsymbol{D}_{\mathrm{id}},\boldsymbol{D}_{\mathrm{expr}}\}$ and the regressors $\{\boldsymbol{R}_{color},\boldsymbol{R}_{gauss}\}$ for the hair and face regions (Fig. 4).
@@ 3.4 Synthetic Dataset
Figure 6: Cross-reenactment comparison of SynShot inversion using only 3 views to state-of-the-art (SOTA) methods: INSTA [85], Flash Avatar (FA) [71], and Splatting Avatar (SA) [59], each of which was trained on an average of 3000 frames. It is evident that without a strong prior, these methods fail to generalize to novel expressions and views. Inversion input images are in the supplemental materials.
@@ 4 Results
For inversion-based methods, we compare PanoHead [2], HeadNeRF [26], and MofaNeRF [84] (see supplemental material for details). We also compare SynShot to newer 4D GAN-based methods offering animation, including Portrait4D [13], Next3D [61], and InvertAvatar [80] (Figures 10 and 8). Moreover, Figure 7 presents a novel view evaluation of challenging long hair and beard inversion, demonstrating strong generalization capabilities of SynShot. We used three input images for all the inversion experiments which can be seen in the supplemental materials.
Evaluation.
To measure the performance of SynShot without introducing bias, we selected training frames from $\{F_{n}\}_{n=1}^{16}=\{1,\dots,987\}$, where $F_{n}$ denotes the Fibonacci sequence. For all experiments, we use progressive farthest point sampling [49] in the 3DMM expression space to select a specified number of frames from the training set. The self-reenactment sequences were evaluated using photometric metrics (LPIPS and SSIM) on the last 600 frames from the INSTA dataset [85].
Refer to captionSourceOursInvertAvatar [80]Portrait4D [13]Next3D [61]
Figure 8: The GAN-based cross-reenactment comparison shows that our method significantly better reconstructs the target subject’s appearance (identity) and remains faithful to the source subject’s head poses and expressions, whereas the other methods suffer from artifacts. Inversion input images can be found in supplemental materials.
@@
Our combined dataset consists of eleven monocular sequences ($512\times 512$ resolution), many of which are in-the-wild videos with very limited head motion, resulting in a low error as the test sequences closely resemble the training data, leaving limited room to assess diversity. To address this and accurately measure the effective error, we trained each method on a varying number of frames, corresponding to frames used in our inversion pipeline. The reconstruction error is evaluated on 600 test frames. Figure 9 demonstrates the effectiveness of our inversion, particularly with up to 233 training frames. Due to the lack of a strong prior, monocular methods fail in low training frame regimes, and, even with larger training datasets, they do not perform well and produce artifacts. Please note that to benefit from an increased number of input frames, i.e., to ground the avatar reconstruction more on the input than the synthetic prior, it requires an increased number of optimization iterations during pivotal fine-tuning. The number of iteration steps affects the reconstruction metrics, causing LPIPS to vary non-monotonically. Refer to the supp. mat. for qualitative evaluation.
@@
Refer to captionGround TruthOursInvertAvatar [80]Portrait4D [13]Next3D [61]
Figure 10: The GAN-based self-reenactment comparison again shows that SynShot better captures identity and synthesizes novel views, proving its usefulness as a synthetic prior and for pivotal fine-tuning in inversion. LPIPS scores: Ours (0.0236), InvertAvatar (0.0962), Portrait4D (0.0843), and Next3D (0.2274). Inversion input images can be found in supplemental materials.
We would like to emphasize the importance of evaluating cross-reenactment, which often reveals issues with generalization and overfitting; however, these aspects are frequently underemphasized, as evaluation sequences are commonly not sufficiently challenging. For instance, Figure 9 indicates that 13 frames may be sufficient for monocular methods to perform well on the test set. Despite achieving high-quality results, most monocular methods [25, 85, 19, 71, 59] struggle with cross-reenactment involving novel expressions and views. In the supp. mat. we present a full evaluation. Without a strong prior, these methods frequently exhibit artifacts when driven by out-of-distribution sequences. In contrast, our method, leveraging only three images and a synthetic prior with effective shape-expression disentanglement, is able to invert an avatar that significantly outperforms state-of-the-art models trained on thousands of frames. Figure 16 demonstrates cross-reenactment, with the leftmost column serving as the source for expression and view. This shows that incorporating a strong prior enhances the visual quality.
@@
GAN-based baselines.
We compared SynShot to three animatable GAN-based methods. For our method and InvertAvatar [80], we used three input images, whereas Portrait4D [13] and Next3D [61] are single-shot. Figure 10 presents qualitative self-reenactment results, with additional quantitative LPIPS scores: Ours (0.0236), InvertAvatar (0.0962), Portrait4D (0.0843), and Next3D (0.2274). Both results show that SynShot significantly outperforms the baselines. Moreover, Figure 8 presents expression transfer, where our method best captures the subject’s identity and is significantly more stable for novel views and expressions, whereas GAN-based methods tend to introduce substantial artifacts, particularly in side views.
VQ-VAE Architecture Ablation.
Loss L1 $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
$\mathcal{L}_{photo}+\mathcal{L}_{VGG}$ 0.0244 0.0839 0.9058 23.1191
Table 2: Ablation for our inversion losses.